# Do holographic CFT states have unique semiclassical bulk duals?

### v1

@@ I Introduction
An important ingredient of these recent calculations is that the inner product is corrected by wormhole contributions. Thus, naively orthogonal semiclassical configurations may be non-orthogonal, at least as CFT states. These wormhole contributions then lead to the factorization problem, i.e. they are in tension with the computation of fine-grained quantities in AdS/CFT unless either a) the bulk theory is dual to an ensemble of boundary theories or b) additional contributions to the path integral exist that cancel their effect. [^1] In option b), wormholes correctly capture the statistics of coarse-grained CFT quantities [9, 10, 11]. The first option is well understood in 2D gravity and by now many examples exist starting from the work of Ref. [12]. However, ensemble averaging of interacting, holographic CFTs in higher dimensions is difficult to understand owing to the sparseness expected of conformal fixed points at finite central charge. The second option is realized by half-wormholes in toy models [13, 14, 15, 16], whereas we currently have no such understanding of the new bulk ingredients in more general theories.
@@ II Closed Universes and Multiple Bulk Descriptions
where heavy means the conformal dimension $\Delta_{\mathcal{O}^{(i)}}$ is above the black hole threshold and $i$ labels different possible choices of matter insertion. [^3] The different choices $i$ need not be related by any symmetry, but could for example be associated with different scaling dimensions. For large $\beta_{L},\beta_{R}$, we expect the state to effectively be below the Hawking-Page transition. By demanding an approximate global symmetry under which the bulk dual of $\mathcal{O}^{(i)}$ is charged, we can ensure that $\mathcal{O}^{(i)}$ has a vanishing one-point function and thus the insertions in the bra and ket must be paired. This leads to a Euclidean saddle of the form shown in Fig. (1). [^4] It would be interesting to improve this saddle by finding smooth solutions supported by smeared matter fields although we see no reason to suspect that the qualitative features would change. The time-symmetric slice suggests that the bulk dual of the state |$\ket{\psi_{i}\left(\beta_{L},\beta_{R}\right)}$ consists of two copies of AdS with radiation that is entangled with a closed universe supported by the matter insertion [34]. We call this Description 1. Note that in Description 1, we have $\braket{\Psi_{i}}{\Psi_{j}}=0$ for $i\neq j$, whereas $|\braket{\Psi_{i}}{\Psi_{j}}|^{2}\sim e^{-S}$ due to a wormhole contribution, where $S$ is the $O(1)$ entanglement entropy of the closed universe with the two copies of AdS [34].
@@
Now we can also use the fact that Eq. (2) has a standard bulk dual using the extrapolate dictionary relating the CFT operator $\Phi$ to the bulk scalar field $\phi$ [39]. This means that we can entirely re-express Eq. (2) as a bulk state involving bulk operators acting on the product of two AdS vacuums. In this description, there is no closed universe and the bulk entanglement is purely between radiation in these two asymptotically AdS universes, see Fig. (2). We call this Description 2. In Description 2, the bulk inner product $\braket{\Psi_{i}}{\Psi_{j}}$ between different states is by construction given by their (non-zero) inner product in the CFT. Given the precise coefficients $c^{(i)}_{jk}$, this bulk description is semiclassical in the sense that it is completely captured by quantum field theory on curved spacetime. Note, however, that the values of the exact coefficients $c_{jk}^{(i)}$ should be regarded as a UV input. In fact, it is difficult to imagine a semiclassical Euclidean saddle being able to compute the fine-grained pseudorandom coefficients, and it instead only captures coarse grained quantities (which behave like averages over the pseudorandom ensemble) via wormholes [9, 10, 11]. This explains the behavior of the inner product in Description 1.
We now have two semiclassical bulk descriptions of the same CFT state and this makes the bulk-boundary dictionary ambiguous. [^7] This ambiguity is of the same type pointed out in Ref. [40] except we have an explicit semiclassical Euclidean saddle constructing the closed universe. Moreover, the state lives in a subspace of dimension $O(1)$, hence its complexity is also at most $O(1)$. Thus, no complexity criteria can resolve the issue. So how can we interpret this result? Below we discuss some possibilities.
@@
Of course, the down side of this option is that we only have an understanding of ensemble averaging in two dimensional theories like JT gravity. In higher dimensional theories, it is difficult to find enough CFTs to ensemble average over. The lack of an ensemble has also been argued on the basis of the swampland cobordism conjecture [41]. On the other hand, some progress towards a definition of a dual ensemble in higher dimensions have been made [18, 42].
@@
Gauge Equivalence
: If we insist on preserving both descriptions in the presence of a single dual CFT, then they must be gauge equivalent in some sense. Since the Marolf-Maxfield Hilbert space of closed universes is one-dimensional, this would require a new framework for describing the semiclassical physics of a closed universe. For instance, since the past and future of the closed universe has a singularity, one could guess that the final state proposal [43] is applied, resulting in a single state. From the point of view of observers in the closed universe, this would modify the usual formalism of quantum mechanics to that of final state quantum mechanics.
This is analogous to the problem of understanding semiclassical physics in cosmology. In the Hartle-Hawking approach [44], there is a single wavefunction of the universe, but it should be possible to make sense of ordinary physics, at least locally. Similarly, in eternal inflation scenarios (see Refs. [45, 46] and Ref. [47] for a review), a single, global wavefunction of the universe should exist, but semiclassical physics should emerge within pocket universes associated with different branches of the wavefunction (see e.g. Ref. [48]). While we have nothing more concrete to say about this option, it is another radical yet interesting option to explore.
AdS $\neq$ CFT
: Finally, if we insist that the closed universe is semiclassical and should be UV completed into some theory of quantum gravity while also using regular quantum mechanics, we would then have to say that AdS/CFT is not a complete description of bulk quantum gravity. There would then be a more complete bulk description which only agrees with AdS/CFT when restricted to questions below the black hole threshold where wormholes are irrelevant [49]. This of course would also be quite drastic, although an interesting such scenario has been proposed recently [50]. Ref. [50] proposed that the bulk theory may potentially violate reflection-positivity in the presence of a large number of asymptotic boundaries without affecting the semiclassical physics of closed universes. Of course, the CFT, being reflection positive, is then unable to capture any such physics.
@@
The following flowchart summarizes the four possibilities outlined above.
@@ III Implications for the Black Hole Interior
Ref. [17] and subsequent papers proposed that the existence of an ensemble makes the calculations of the Page curve consistent. Of course, one may then ask what happens if we calculate the Page curve in an $\alpha$ state. In this case, if we extend the analogy from the closed universe, the black hole interior will be non-semiclassical, i.e. it stops existing. This also continues to be the case if we use option b).
Option c) is of course familiar in the literature as the final state proposal of Ref. [43]. As has been discussed earlier, there are various issues with this proposal as well [53]. Option d) on the other hand says that the black hole interior potentially exists, but AdS/CFT cannot tell us anything about it.
@@ IV Implications for the gravitational path integral
Finally, we would like to comment briefly about the implications of the four options outlined above for the role of the path integral in quantum gravity.
@@
If option b) is correct, then the path integral, at least in the saddle point approximation, is only able to capture coarse-grained quantities, which behave like averages over an ensemble due to the pseudorandom statistics of the coefficients $c_{jk}^{(i)}$ defined in equation (3) [9, 10, 11]. An interesting question is whether the full, non-perturbative path integral is able to compute fine-grained quantities, which should factorize, and yield Description 2. This was shown to be possible in toy models, where half-wormholes rescue factorization [13, 14, 15, 16]. It seems plausible that additional ingredients (such as strings, branes, and other UV objects) need to be included in the path integral in order to obtain this result, as it was the case in Refs. [13, 14, 15, 16].
Drawing conclusions for the path integral in options c) and d) is harder because of their radical, unconventional nature. Option c) suggests that the path integral is able to compute CFT quantities such as overlaps between different states in the same way as option b), but it might not be able to capture the new framework needed to describe semiclassical physics emerging within the one-dimensional Hilbert space of the closed universe.
Finally, option d) could imply that the gravitational path integral, at least the way it is usually defined, consistently computes bulk observables—in a similar way to option b)—only when the bulk quantum gravity has a dual description in terms of a CFT. When it does not, a different definition of the path integral might be needed (e.g. one violating reflection positivity as proposed in [50]), or one might not exist at all. This would not be shocking, given the well-known difficulties in defining the bulk quantum gravity path integral independently of a dual boundary path integral (see [54, 55] and references therein).
@@
Acknowledgments: We are thankful to Raphael Bousso, Xi Dong, Thomas Hartman, Adam Levine, Javier Magan, Don Marolf, Henry Maxfield, Shiraz Minwalla, Geoff Penington, Mukund Rangamani, Phil Saad, Martin Sasieta, Arvin Shahbazi-Moghaddam, Brian Swingle, and Herman Verlinde for useful conversations related to the topics of the essay. This work was supported in part by the Berkeley Center for Theoretical Physics; and by the Department of Energy, Office of Science, Office of High Energy Physics under QuantISED Award DE-SC0019380. S.A. is supported by the U.S. Department of Energy through DE-FOA-0002563.

### v3

@@ I Introduction
An important ingredient of these recent calculations is that the inner product is corrected by wormhole contributions. Thus, naively orthogonal semiclassical configurations may be non-orthogonal, at least as CFT states. These wormhole contributions then lead to the factorization problem, i.e. they are in tension with the computation of fine-grained quantities in AdS/CFT unless either a) the bulk theory is dual to an ensemble of boundary theories or b) additional contributions to the path integral exist that cancel their effect. [^1] In option b), wormholes correctly capture the statistics of coarse-grained CFT quantities [9, 10, 11, 12]. The first option is well understood in 2D gravity and by now many examples exist starting from the work of Ref. [13]. However, ensemble averaging of interacting, holographic CFTs in higher dimensions is difficult to understand owing to the sparseness expected of conformal fixed points at finite central charge. The second option is realized by half-wormholes in toy models [14, 15, 16, 17], whereas we currently have no such understanding of the new bulk ingredients in more general theories.
@@ II Closed Universes and Multiple Bulk Descriptions
where heavy means the conformal dimension $\Delta_{\mathcal{O}^{(i)}}$ is above the black hole threshold and $i$ labels different possible choices of matter insertion. [^3] The different choices $i$ need not be related by any symmetry, but could for example be associated with different scaling dimensions. For large $\beta_{L},\beta_{R}$, we expect the state to effectively be below the Hawking-Page transition. By demanding an approximate global symmetry under which the bulk dual of $\mathcal{O}^{(i)}$ is charged, we can ensure that $\mathcal{O}^{(i)}$ has a vanishing one-point function and thus the insertions in the bra and ket must be paired. This leads to the dominant Euclidean saddle being of the form shown in Fig. (1). [^4] It would be interesting to improve this saddle by finding smooth solutions supported by smeared matter fields although we see no reason to suspect that the qualitative features would change. The time-symmetric slice suggests that the bulk dual of the state |$\ket{\psi_{i}\left(\beta_{L},\beta_{R}\right)}$ consists of two copies of AdS with radiation that is entangled with a closed universe supported by the matter insertion [35]. We call this Description 1. Note that in Description 1, we have $\braket{\Psi_{i}}{\Psi_{j}}=0$ for $i\neq j$, whereas $|\braket{\Psi_{i}}{\Psi_{j}}|^{2}\sim e^{-S}$ due to a wormhole contribution, where $S$ is the $O(1)$ entanglement entropy of the closed universe with the two copies of AdS [35].
@@
Now we can also use the fact that Eq. (2) has a standard bulk dual using the extrapolate dictionary relating the CFT operator $\Phi$ to the bulk scalar field $\phi$ [40]. This means that we can entirely re-express Eq. (2) as a bulk state involving bulk operators acting on the product of two AdS vacuums. In this description, there is no closed universe and the bulk entanglement is purely between radiation in these two asymptotically AdS universes, see Fig. (2). We call this Description 2. In Description 2, the bulk inner product $\braket{\Psi_{i}}{\Psi_{j}}$ between different states is by construction given by their (non-zero) inner product in the CFT. Given the precise coefficients $c^{(i)}_{jk}$, this bulk description is semiclassical in the sense that it is completely captured by quantum field theory on curved spacetime. Note, however, that the values of the exact coefficients $c_{jk}^{(i)}$ should be regarded as a UV input. In fact, it is difficult to imagine a semiclassical Euclidean saddle being able to compute the fine-grained pseudorandom coefficients, and it instead only captures coarse grained quantities (which behave like averages over the pseudorandom ensemble) via wormholes [9, 10, 11, 12]. This explains the behavior of the inner product in Description 1.
We now have two semiclassical bulk descriptions of the same CFT state and this makes the bulk-boundary dictionary ambiguous. [^7] This ambiguity is of the same type pointed out in Ref. [41] except we have an explicit semiclassical Euclidean saddle constructing the closed universe. A similar puzzle in the evaporating black hole context was also pointed out and discussed in [42, 43]. Moreover, the state lives in a subspace of dimension $O(1)$, hence its complexity is also at most $O(1)$. Thus, no complexity criteria can resolve the issue. So how can we interpret this result? Below we discuss some possibilities.
@@
Of course, the down side of this option is that we only have an understanding of ensemble averaging in two dimensional theories like JT gravity. In higher dimensional theories, it is difficult to find enough CFTs to ensemble average over. The lack of an ensemble has also been argued on the basis of the swampland cobordism conjecture [44]. On the other hand, some progress towards a definition of a dual ensemble in higher dimensions have been made [19, 45, 46].
@@
Beyond AdS/CFT
: If we assume that a bulk observer in a closed universe should experience ordinary semiclassical physics [^8] This is a well-motivated assumption: there is no conclusive evidence that we do not live in a closed universe [47, 48, 49], and our local experience should be independent of the global properties of the universe. and that it should be possible to describe a closed universe in a UV-complete theory of quantum gravity, we would have to conclude that AdS/CFT is not a complete description of bulk quantum gravity. There would then be a more complete bulk description able to capture an observer’s experience, which only agrees with AdS/CFT when restricted to global questions asked from a CFT point of view. This, of course, would also be quite drastic, although interesting such scenarios have been proposed recently [50, 51]. Ref. [50] proposed that the bulk theory may potentially violate reflection-positivity in the presence of a large number of asymptotic boundaries without affecting the semiclassical physics of closed universes. Of course, the CFT, being reflection positive, is then unable to capture any such physics. Ref. [51], on the other hand, proposed a new way to describe quantum gravity in the presence of an observer using the gravitational path integral, and concluded that the Hilbert space relevant to describe the observer’s experience is much larger than the Hilbert space of the CFT living on the asymptotic AdS boundary.
@@
The following flowchart summarizes the three possibilities outlined above.
@@ III Implications for the Black Hole Interior
Ref. [18] and, with a slightly different approach, Refs. [42, 43] proposed that the existence of an ensemble makes the calculations of the Page curve consistent. Of course, one may then ask what happens if we calculate the Page curve in an $\alpha$ state. In this case, if we extend the analogy from the closed universe, the black hole interior will be non-semiclassical, i.e. it stops existing. This also continues to be the case if we use option b). A bulk picture compatible with option b) is also provided by the final state proposal of Ref. [54], in which semiclassical physics in the black hole interior breaks down (see [55] for a discussion of various issues with this proposal). Option c) on the other hand says that the black hole interior potentially exists, but AdS/CFT cannot tell us anything about it.
@@ IV Implications for the gravitational path integral
Finally, we would like to comment briefly about the implications of the three options outlined above for the role of the path integral in quantum gravity.
@@
If option b) is correct, then the path integral, at least in the saddle point approximation, is only able to capture coarse-grained quantities, which behave like averages over an ensemble due to the pseudorandom statistics of the coefficients $c_{jk}^{(i)}$ defined in equation (3) [9, 10, 11, 12]. An interesting question is whether the full, non-perturbative path integral is able to compute fine-grained quantities, which should factorize, and yield Description 2. This was shown to be possible in toy models, where half-wormholes rescue factorization [14, 15, 16, 17]. It seems plausible that additional ingredients (such as strings, branes, and other UV objects) need to be included in the path integral in order to obtain this result, as it was the case in Refs. [14, 15, 16, 17], see Figure 3.
Figure 3: Pictorial representation of how the non-perturbative gravitational path integral could yield Description 2. When additional, non-perturbative ingredients are taken into account, “half-wormhole”-like saddles (center figure) can contribute to the path integral [14, 15, 16, 17], with the closed universe replaced by some non-semiclassical correlation. The sum of such saddles and the semiclassical saddle associated with Description 1 (left figure) could be equivalent to a Euclidean geometry without a closed universe (right figure), which would lead to Description 2.
Finally, option c) could imply that the gravitational path integral, at least the way it is usually defined, consistently computes bulk observables only when they correspond to well-defined observables in the boundary CFT. In particular, such observables should act linearly on the CFT Hilbert space. When they do not—as it is the case, e.g., for the experience of an observer in a closed universe or infalling into an old black hole, when observables become state-dependent and therefore non-linear on the CFT Hilbert space—a different definition of the path integral might be needed. For instance, this new definition could involve taking into account the presence of an observer as proposed in [51], or violating reflection positivity as proposed in [50], or it might not exist at all. This would not be shocking, given the well-known difficulties in defining the bulk quantum gravity path integral in higher dimensions independently of a dual boundary path integral (see [56, 57] and references therein).
@@
Acknowledgments: We are thankful to Alex Belin, Raphael Bousso, Xi Dong, Daniel Harlow, Thomas Hartman, Adam Levine, Javier Magan, Don Marolf, Henry Maxfield, Shiraz Minwalla, Geoff Penington, Mukund Rangamani, Phil Saad, Martin Sasieta, Arvin Shahbazi-Moghaddam, Brian Swingle, and Herman Verlinde for useful conversations related to the topics of the essay. This work was supported in part by the Berkeley Center for Theoretical Physics; and by the Department of Energy, Office of Science, Office of High Energy Physics under QuantISED Award DE-SC0019380. S.A. is supported by the U.S. Department of Energy through DE-FOA-0002563.